OpenStudy (anonymous):

Find equations of all lines tangent to y=9-x^2 that pass through the point (1,12).

hero (hero):

This can be done without calculus right? @satellite73

OpenStudy (anonymous):

i don't think so

hero (hero):

Okay @iloveme17. You seem to already know how to do it. What is the difficulty here?

OpenStudy (anonymous):

how do i seem to know how to do it i have no idea how to ...

OpenStudy (anonymous):

ok, you know the derivative is \(-2x\) right? and any point on the curve looks like \((x,9-x^2)\) the slope between that general point and the specific point \((1,12)\) will be \[m=\frac{9-x^2-12}{x-1}\] and that must be equal to \(-2x\)

OpenStudy (anonymous):

so your job is to set \[\frac{9-x^2-12}{x-1}=-2x\] \[\frac{-3-x^2}{x-1}=-2x\] etc it is a quadratic , that is why you get two answers

OpenStudy (anonymous):

@hero i am not sure of how you can do this without using calc, mainly because the notion of a tangent line is a calc idea. that is not to say there isn't some other way

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw i take it you are good from there